OpenStudy (anonymous):

evulate x^2+3y when x=-0.1 and y=-2

OpenStudy (ksaimouli):



OpenStudy (anonymous):

what does that mean?

OpenStudy (unklerhaukus):

\[x^2+3y\] when \(x=-0.1\) and \(y=-2\) \[(-0.1)^2+3(-2)\]\[=-0.1\times-0.1+3\times-2\]

OpenStudy (anonymous):

its either -5.99 or 5.99 or 6.1 or -6.01

OpenStudy (unklerhaukus):

\[=(−0.1×−0.1)+(3×−2)\] \[=0.01-6\]

OpenStudy (anonymous):

thanks